A company has just paid a dividend of 3.28$. Its discount rate is 8.9%, and the expected perpetual growth rate is 5.5%. What would you expect to be the stock's price TODAY?

Express your answer in dollars, rounded to the nearest cent (2 decimals).
Answer #1

Stock price = next year divided /(discount rate - growth rate)

Next year divided = current divided *(1+ growth rate)

=3.28*(1+0.055 )= 3.4604

Stock price = 3.4604/(8.9%- 5.5%)

= 3.4604/3.40%

= 101.78